Let G be a finite group. Show that the number of elements g of G such that g3= e is

odd, where e is the identity of G.


1
Expert's answer
2021-06-10T06:59:52-0400

Let "G" be a finite group. Let us show that the number of elements "g\in G" such that "g^3= e" is

odd, where "e" is the identity of "G". If "g^3=e", then "(g^{-1})^3=g^{-3}=(g^3)^{-1}=e^{-1}=e". Let us find all "g\in G" such that "g^3=e" and "g^{-1}=g." In this case "g^2=e," and hence "g=ge=gg^2=g^3=e."  Therefore, for all non-identity elements "g" with "g^3=e" we have that "(g^{-1})^3=e" and the set "\{g,g^{-1}\}" is of cardinality 2. Since "e^3=e" and "|\{g,g^{-1}\}|=2" for "g\in G\setminus\{e\}" and "g^3=e", we conclude that the number of elements "g\in G" such that "g^3= e" is equal to  "|\{e\}\cup(\bigcup\limits_{g\in G\setminus\{e\},\ g^3=e}\{g,g^{-1}\})|=1+2k" (where "k" is a number of different pairs "\{g,g^{-1}\}"), and hence is odd.